OpenStudy (anonymous):

3x^4+5x^3+25x^2+45x-18 zeros:3i,-3i. I need help finding the other two zeros

OpenStudy (anonymous):

if those are the two zeros, then one factor is \((x+3i)(x+3i)=x^2+9\)

OpenStudy (anonymous):

Yeah I got that part

OpenStudy (anonymous):

that means \[3x^4+5x^3+25x^2+45x-18 =(x^2+9)(\text{something})\] find the something by either thinking or by division

OpenStudy (anonymous):

i like the think method, it is easier than dividing

OpenStudy (anonymous):

for example, the first term has to be \(3x^2\) otherwise you are not going to get \(3x^4\) when you multiply

OpenStudy (anonymous):

the last term has to be \(-2\) so you will get \(-18\) when you multiply

OpenStudy (anonymous):

\[3x^4+5x^3+25x^2+45x-18 =(x^2+9)(3x^2+bx-2)\] and all you need is \(b\)

OpenStudy (anonymous):

Is the b 5x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[3x^2+5x-2\] is the other factor, which in turn factors as \[(3x-1)(x+2)\]

OpenStudy (anonymous):

OMG thank you

OpenStudy (anonymous):

yw